Shift summary for day shift, 26/Nov/99

AG (ag@fnal.gov)
Fri, 26 Nov 1999 16:27:45 -0500 (EST)

Shift Summary for day shift, 26/Nov/99

On Shift: MA, AG

Hours of beam available at all intensities: 6.5h
Hours of beam available at nominal intensity: 6.0h
Hours of E799 data (special runs not included): 5.2h
Hours of special runs (laser scans, muon runs...): 0.0h

Accounting of lost time:

o errors while starting run 15059
o tape change, dart fresh start
o booster extraction problems
o run 15063 abort
o another tape change, dart fresh start again

========================================================================

Runs ended this shift:
Run 15062 - < epsilon - BEAM_799
Run 15063 - 0.3M - BEAM_799

Tapes brought to FCC this shift:
(including tapes for on-line split)

UP[A-C]18[4-9] and UPS06[2-3],
UP[A-C]19[3-5] and UPS065.

========================================================================

Accesses during shift:

none

========================================================================

Overview of shift:

Rather problematic E799 data-taking in the first half of the shift. First,
failed to start run 15059 normally after tapes filled up ( by the end of
the owl shift ) and had to clean tape drives and replace new set of tapes
without writing any data on them. Then lost beam due to booster extraction
problems. Later in the shift run 15063 aborted and ended with tape error.
After another tape change and fresh start continued to take data through
the rest of the shift.

========================================================================

Description of Shift:

run spill time comment

------------------------------------------------------------------------

15059_________0700_______Day shift begins. Run 15059 is being started
_________________________(with fresh tapes 190-192, S064) after tapes
_________________________filled up earlier in the owl shift.

______________0710_______"ocprc da_start: Error starting run 15059"
_________________________try to fresh start -> can't even initialize
_________________________try to logout from the da_ktev3 session, then
_________________________do fresh start again -> still can't pass even
_________________________"stage 1" of initialization...
_________________________Paged Leo -> Viv ... Leo is coming over ...
_________________________Meanwhile all tape drives come up red in the
_________________________initialization status window -> decide to clean
_________________________all the tape drives and put new set of tapes ...

_________________________Viv traced the problem to the bad tape (tape
_________________________drive ?) - ktev4,drv4.

15062_________0755_______Now did fresh start and it works this time, but
_________________________beam is gone due to booster extraction problems.

______________0907_______Still no beam. Ed calls and asks to stop/start
_________________________the run in order to pick up new CsI constants
_________________________(that should fix ch 1096 problem). He added an
_________________________extra gain boost to 1096 above what calibration
_________________________gave.

15063_________0920_______Run 15063 started. Booster problems continue -
_________________________experts en-route, no estimate.

______________1016_______One spill of beam at 1e12.

______________1019_______Beam at 2e12, they are tuning it back on...

______________1118_______Run aborts and stops with tape error. Don't know
_________________________about EOF marks, so change tapes and do fresh
_________________________start (after a helpful conversation with Sydney
_________________________and voluntary support from JasonL; see details
_________________________in the KTeV LogXXII, p.2).

15064_________1132_______Got another run started...

______________1340_______Mounted UPH034,UPF049 for ONSPLIT.

========================================================================

Follow up Notes on Previous Shift

========================================================================

Items to Follow up on Next Shift

Channels 1476,2893 had gain drift of ~ -5% during this shift. Continue
monitoring...